# IceCube Collaboration

### v1

@@ Appendix B Parameter impact on analysis sample
In our parametrization, the GENIE+HKKMS’15 model is best described by $\langle y\rangle$ $=0.43\pm 0.02$ and $\log_{10}\lambda=0.00^{+0.11}_{-0.08}$ in the energy range of interest. In order to illustrate the impact of assuming different values of $\langle y\rangle$ and $\lambda$, we calculate the expected change in number of events in the analysis binning of reconstructed energy and inelasticity. Figure 15 shows three example $\frac{dp}{dy}$ distributions with different values of $\langle y\rangle$ and $\lambda$. The comparison between the event counts corresponding to all three distributions and the expectation from GENIE+HKKMS’15 is shown on Figure 16. Figure 17 shows changes in the number of events due to modifications in systematic parameter values, while the GENIE+HKKMS’15 model is assumed for the inelasticity distribution.
@@
In order to quantify correlations between physics and nuisance parameters in the fit, we performed a pseudo-trial test. We generated an ensemble of 400 pseudo-experiments assuming nominal values of fit parameters, but including Poisson fluctuations in each analysis bin. For each of the pseudo-experiments an inelasticity distribution fit was performed and resulting posterior parameter distributions were used to calculate Pearson correlation coefficients for each pair of fit parameters. Figure 18 shows the Pearson correlation matrix based on all 400 trials in the test.
Figure 18: Parameter correlations based on ensemble test at the best fit point.
After the fit to data was done, we also conducted a two-dimensional likelihood scan for all 15 unique pairs of physics parameters. Figure 19 shows two-dimensional likelihood scans for each pair of parameters, featuring best fit point and confidence level contours.
Figure 19: Two-dimensional likelihood scans for all unique combination of physics parameters. Black cross corresponds to the best fit point and the contours show 1, 2 and 3 $\sigma$ confidence intervals.

### v2

@@ Appendix B Parameter impact on analysis sample
In our parametrization, the genie+HKKMS’15 model is best described by $\langle y\rangle$ $=0.43\pm 0.02$ and $\log_{10}\lambda=0.00^{+0.11}_{-0.08}$ in the energy range of interest. To illustrate the impact of assuming different values of $\langle y\rangle$ and $\lambda$, we calculate the expected change in number of events in the analysis binning of reconstructed energy and inelasticity. Figure 15 shows three example $\frac{dp}{dy}$ distributions with different values of $\langle y\rangle$ and $\lambda$. The comparison between the event counts corresponding to all three distributions and the expectation from genie+HKKMS’15 is shown on Fig. 16. Figure 17 shows changes in the number of events due to modifications in systematic parameter values, while the genie+HKKMS’15 model is assumed for the inelasticity distribution.
@@
To quantify correlations between physics and nuisance parameters in the fit, we performed a pseudotrial test. We generated an ensemble of 400 pseudoexperiments assuming nominal values of fit parameters, but including Poisson fluctuations in each analysis bin. For each of the pseudoexperiments an inelasticity distribution fit was performed, and resulting posterior parameter distributions were used to calculate Pearson correlation coefficients for each pair of fit parameters. Figure 18 shows the Pearson correlation matrix based on all 400 trials in the test.
Figure 18: Parameter correlations based on ensemble test at the best-fit point.
After the fit to data were done, we also conducted a two-dimensional likelihood scan for all 15 unique pairs of physics parameters. Figure 19 shows two-dimensional likelihood scans for each pair of parameters, featuring best-fit point and confidence level contours.
[…]